OpenStudy (smurfy14):

Is this and even, odd, or neither function: f(x)=x^3+2

OpenStudy (anonymous):

Even function - symmetrical about the y-axis. Odd function - symmetrical about the origin. Which one does f(x) fall into? Graph it. :)

OpenStudy (gg):

if f(-x)=f(x) it's even , if f(-x)=-f(x) it's odd function. now u just have to check it f(-x)= (-x)^3+2 = (-x)(-x)(-x) +2 = -x^3 + 2 this is neither even nor odd

OpenStudy (anonymous):

Yep, that's the notational form. Proceed

OpenStudy (smurfy14):

thanks you guys!

OpenStudy (gg):

I'm girl ;)

OpenStudy (smurfy14):

lol sry